OpenStudy (snickers):

x^5+x^2/x^-2 simplifiy

OpenStudy (anonymous):

negative exponents are reciprocals, so if its negative exponent just put it on the other side of the fraction: x^5 + x^2/x^-2 = (x^5+x^2)(x^2)= x^7+x^4

OpenStudy (anonymous):

another way to look at it, is that when you divide exponents with the same bases (base is x in this case) you subtract the exponents value from the numerator: x^5+x^2/x^-2= x^(5-[-2])+x^(2-[-2])= x^(5+2)+x^(2+2)= x^7+x^4

OpenStudy (anonymous):

so theres: \[x^m \div x^{-n} = x^m \times x^{n} \] and: \[x^m \div x^{n} = x^{m-n}\]

OpenStudy (anonymous):

=x^5+x^2+2 =x^5+x^4 =x^4(x+1)